# Bayesian inference for the Markov-modulated Poisson process with an outcome process

### v1

@@ Abstract
In medical research, understanding changes in outcome measurements is crucial for inferring shifts in a patient’s underlying health condition. While data from clinical and administrative systems hold promise for advancing this understanding, traditional methods for modelling disease progression struggle with analyzing a large volume of longitudinal data collected irregularly and do not account for the phenomenon where the poorer an individual’s health, the more frequently they interact with the healthcare system. In addition, data from the claim and health care system provide no information for terminating events, such as death. To address these challenges, we start from the continuous-time hidden Markov model (CTHMM) that has been used previously to understand disease progression by modelling the observed data as an outcome whose distribution depends on the state of a latent Markov chain representing the underlying health state. However, we also allow the underlying health state to influence the timings of the observations via a point process. Furthermore, we create an addition “death” state and model the unobserved terminating event, a transition to this state, via an additional Poisson process whose rate depends on the latent state of the Markov chain. This extension allows us to model disease severity and death not only based on the types of care received but also on the temporal and frequency aspects of different observed events. We present an exact Gibbs sampler procedure that alternates sampling the complete path of the hidden chain (the latent health state throughout the observation window) for each individual conditional on the parameters with sampling the parameters conditional on the complete paths. When the unobserved, terminating event occurs early in the observation window, there are no more observed events, and naive use of a model with only “live” health states would lead to biases in parameter estimates; our inclusion of a “death” state mitigates against this. Simulation studies demonstrate that the proposed Gibbs sampler performs effectively. Finally, we apply our method to health care claim data from a Canadian cohort, obtaining valuable insights for healthcare management.
@@ 1 Introduction
A multi-state model captures the status of an individual longitudinally as a time-discretization of a continuous-time Markov process, and in many settings the common assumption that the health status of a patient falls into one of a finite number of categories (usually ordered by severity) is a plausible one. Such a process can be modeled as a continuous-time hidden Markov model (CTHMM), which has been studied extensively in medical applications such as disease progression (Luo et al., 2021) and cancer screening (Lange et al., 2015; Klausch et al., 2023). While CTHMM can address the issue of non-equidistant longitudinal observations, in medical research, the observation times themselves might be a vital sign of the illness of a patient, and it is possible that the timing of healthcare utilization is informative in the assessment of health status progression. For example, patients who are in more severe health states typically interact with the healthcare system more frequently. There have been studies modelling the observation process itself using frequentist approaches, such as Lange et al. (2015) and Mews et al. (2023). In addition, data from the claim and health care system provide no information for terminating events, such as death. For example, patients with COPD over the age of 65 are at high risk of death. Modelling death would provide useful information for how well they are managed in the system. Nevala et al. (2024) included the death data from a different data system to model the colorectal cancer progression. However, from our knowledge, there is no existing method to account for the terminating event, such as death, when it is not available in the dataset. In this paper, we develop a Bayesian inference procedure that takes into account the observation process by modelling the observed data as an outcome process and the time of observations as a point process jointly alongside the CTHMM, given the underlying latent health state. Moreover, we also consider the death as missing data, and model the death as an informative censoring. This extension allows us to model disease severity and death not only from the types of care received but also from the times between different observed events and their respective frequency. It also provides insights on how long patients tend to survive in each state.
@@
The paper is organized as follows. Section 2 demonstrate the model structure, followed by the likelihood construction in Section 2.1 and the forward-backward method to calculate the probability of each state at the observation time in Section 3. Section 4 layouts the Gibbs sampler to draw inference for the model, and some empirical studies, including simulation and real data analysis, are in Section 5. The paper concludes in Section 6 with a summary of findings and outlines areas for future research.
@@ 2 A Markov-modulated Poisson process with an outcome process
Conditional on the latent process being in state $k\in\mathcal{K}$, at time $\tau_{t}$, the outcome observation $o_{t}$ is drawn from a distribution with a probability mass or density function $f_{t}(o_{t}|X_{\tau_{t}}=k)\equiv f_{k}(o_{t}|z_{t},\beta_{k})$, where $z_{t}\in\mathbb{R}^{D}$ is a vector of covariates (including an intercept) and $\beta$ is a $D$-vector of covariate effects. In all of our simulations, including the real-data example, the prior for $\beta_{k}$, is chosen to be conditionally conjugate to $f_{k}$ so that it is straightforward to sample from the conditional posterior for each $\beta_{k}$ given $X_{\tau_{1}},\dots,X_{\tau_{T}}$. Non-conjugate case can easily be extended and sampled using the Metropolis-Hastings algorithm. Define matrix $\mathbf{B}=(\beta_{d,k})$ for $d=1,\ldots,D$ and $k=1,\ldots,K$ as the regression coefficient matrix.
@@ 2.1 Likelihood for the MMPP with an outcome process
$\mathcal{L}(\mathbf{Q},\nu|x_{[0,\tau]})=f(x_{[0,\tau]}|\mathbf{Q},\nu)=\nu_{x_{0}}\prod\limits_{l=1}^{K}{\prod\limits_{m\neq l}{{q_{l,m}}^{{N_{l,m}}}\exp\left({-{q_{l,m}}{R_{l}}}\right)}}$
@@ 3 The forward-backward algorithm
$\left(\begin{array}[]{cc}Q-\Lambda&\boldsymbol{\lambda}^{\top}\\ \mathbf{0}^{\top}&0\\ \end{array}\right)$
where $\Lambda=\text{diag}\left(\lambda_{1},\ldots,\lambda_{K}\right)$. Writing $Y_{\tau^{-}}$ for $\lim_{s\uparrow\tau}Y_{s}$, and defining $\eta_{t-1,t,i,k}:=\left[\exp\left\{(Q-\Lambda)(\tau_{t}-\tau_{t-1})\right\}\right]_{i,k}$, we see that
$\mathbb{P}\left(\tau_{t},X_{\tau_{t}}=k\left|X_{{\tau_{t-1}}}=i\right.\right)=\mathbb{P}\left(Y_{\tau_{t}^{-}}=k\left|Y_{\tau_{t-1}}=i\right.\right)\lambda_{k}=\left[\exp\left\{(Q-\Lambda)(\tau_{t}-\tau_{t-1})\right\}\right]_{i,k}\lambda_{k}=\eta_{t-1,t,i,k}\lambda_{k}.$
@@
$\mathbb{P}(X_{\tau_{e}}=k,\tau_{e}|X_{\tau_{T}}=i)=\left[\exp\left\{(Q-\Lambda)(\tau_{e}-\tau_{T})\right\}\right]_{i,k}=:\eta_{T,e,i,k}.$
@@
From its definition, $a_{e,k}\propto\mathbb{P}(X_{\tau_{e}}=k|o_{1:T},\tau_{1:T},\tau_{e})$ so we may sample a value for $x_{\tau_{e}}$ with the probabilities of the individual states proportional to the elements of $a_{e}$. Now we suppose that we have simulated $x_{\tau_{e}},x_{\tau_{T}},\dots,x_{\tau_{t+1}}$ from their joint distribution given $o_{1:T}$, $\tau_{1:T}$ and $\tau_{e}$.
We now describe how to simulate $x_{\tau_{t}}$ backward from its conditional distribution given $x_{\tau_{e}},x_{\tau_{T}},\dots,x_{\tau_{t+1}}$, $o_{1:T}$, $\tau_{1:T}$ and $\tau_{e}$. From the hidden-Markov structure, conditional on $x_{\tau_{t+1}}$, all random variables with time indices before $t+1$ are independent of those with indices after $t+1$, so
@@ 5.1 Example 1
For each of $N=50$ subjects, we simulate a realisation of a two-state latent process without the death, $X_{[0,5]}$, using a generator of
@@
Trace plots, auto-correlations plots and histograms for $\lambda_{1}$ and $q_{12}$ are shown in Figures 3 and 4; Integrated auto correlation times (IACTs) for $\lambda_{1}$, $\lambda_{2}$, $q_{1,2}$, $q_{2,1}$, $\beta_{1}$ and $\beta_{2}$ are provided in Table 1. In Scenario I, where the data are generated from a model characterized by clear separation between the states, the Gibbs sampler performs well, with fast convergence and mixing. In Scenario II, where the timing of the events provides no information about the state, the Gibbs sampler’s performance declines slightly, though it still remains promising. The posterior distribution of $\lambda_{1}$ is slightly off the true value. Challenges arise when the data are generated from a less well-separated outcome process, leading to notable auto-correlations in the posterior samples, particularly those associated with $\mathbf{Q}$. The reduction in mixing efficiency is considerable; however, the chain does still mix adequately over the 10,000 iterations shown. By contrast, in Scenario IV, when the point-process information is removed from Scenario III, the exploration of the posterior over the 10,000 iterations is so poor that we would judge that the chain has probably not yet even converged. Moreover, the posterior for $q_{1,2}$ includes values that are not strongly supported when all the data are used. This illustrates the most important point: especially in cases where the outcome process is not particularly informative, the information from the timings of the events provides an important contribution to the posterior and, possibly because of the tighter posterior and more clearly defined paths for each $X_{[0,\tau]}$, improves the mixing of the chain.
@@ 5.2 Example 2
In this example, we demonstrate valid inference for simulated data with a similar pattern to our real data. We generate the data from a three-state model which is similar to that used for our real data, with $N=500$ subjects. For each subject, we simulate a realisation of a three-state latent process with the death as the absorbing state, $X_{[0,10]}$,
@@
This is equivalent to $O_{t}|(Z_{t}=1)\sim\mathsf{Poisson}(\mu_{1,k})$ and $O_{t}|(Z_{t}=0)\sim\mathsf{Poisson}(\mu_{0,k})$, with $\log\mu_{1,k}=B_{1,k}+B_{2,k}$ and $\log\mu_{0,k}=B_{1,k}$. In this case, the outcome expectations (marginalised over the covariate) are $0.35\mu_{0,k}+0.65\mu_{1,k}$, which are $0.46$ and $1.71$ for the two states, respectively. The point process rate is set as $\lambda=(6,10,0)^{\top}$. State 3 is the death state without any observed outcome. If the patient progresses to the absorbing state before the observation window, i.e. death, then the last observation is the final one before death. Otherwise, the last observation is the one right before the observation window, $\tau=10$. We assumed that the first observation is made at time 0 according to the initial distribution. Similar to Example 1, we place non-informative priors on all the parameters in this study. Each of $\lambda_{1}$, $\lambda_{2}$, $q_{12}$ and $q_{21}$ have independent $\mathsf{Gamma}(1,1/8)$ priors. We place independent $\mathsf{Gamma}(0.1,0.1)$ priors on each of the four Poisson expectations, $\mu_{i,k}$, $i=0,1$, $k=1,2$, and a $\mathsf{Beta}(1,1)$ prior for $\nu_{1}$.
Table 2 presents the integrated auto correlation times (IACTs) for $\boldsymbol{\lambda}$, $\mathbf{Q}$ and $\mathbf{B}$ and shows that all of these parameters mix well. Figure 5 presents a histogram of the marginal posterior for each of these parameters, with each true parameter value falling within the main posterior mass. This example demonstrate that by simulating the data-induced window similar to the real data, we can achieve valid inference.
@@ 5.3 Application: Canadian healthcare system data
In this section, we delve into the proposed methodology designed to enhance our comprehension of the progression of chronic diseases, particularly COPD. The significance of this understanding lies in its potential to facilitate early diagnosis, personalized care, and effective management of healthcare systems. While clinical and administrative data harbor valuable insights, traditional methods for modelling disease progression face limitations when dealing with data collected at irregular intervals and interaction frequency, such as when patients choose to engage with the healthcare system.
@@
Utilizing established case definitions based on diagnostic codes (Lix et al., 2018), a total of 76,888 COPD patients were enrolled, with an incident event (ICD-9 491x, 492x, 496x; ICD-10 J41-J44) occurring after a minimum of two years at risk with no events. The observation period spanned from January 1998, commencing with the patients’ first diagnosis, until December 2014. Physicians observed these patients solely during medical visits, which transpired when patients chose to interact with the healthcare system. In these medical visits, pertinent medical information was gathered, and from this observed information, we aimed to infer the patients’ unobserved disease status, where is represented using a discrete disease state model. In this study, their disease status was indirectly gauged through a proxy: the number of prescribed medications at the time of observation. Notably, this information was available exclusively for patients with drug insurance, leading us to focus on patients aged over 65 years, constituting the study population for our analysis. This full dataset has been previously analysed in Luo et al. (2021). Due to the data restriction, we are allowed to randomly sample 1,000 patients to demonstrate our proposed method. Figure 6 shows one COPD patient’s follow up from this dataset. As shown in the figure, we observe the number drugs prescribed as the observation process. In addition, we have also observed variations in the type of encounters, i.e., general practitioner visit (GP), hospitalization (HOSP), specialist visit (SPEC) and emergency department visit (ED), which will be incorporated as time-varying covariates in our model. These changes in the observation point process, transitioning from sparse to dense, reflect the evolving severity of the chronic disease under study.
@@
We apply the model from Section 2 to this sample dataset, with the inference methodology from Section 4. We focus on a three-state model with State 3 (the unobserved death state) as an absorbing state. We place independent $\mathsf{Gamma}(1,1/8)$ priors on each $\lambda_{i}$ and all elements in $\mathbf{Q}$, and a $\mathsf{Gamma}(0.1,0.1)$ prior on each expected number of drugs dispensed for each healthcare utilization, and a $\mathsf{Beta}(1,1)$ prior for $(\nu_{1},\nu_{2})$, with $\nu_{3}=0$. The outcome process is modelled with a Poisson distribution, with the types of health care utilization (HOSP, SPEC, GP, and ED) as a time-dependent covariate; equivalently, for each of states 1 and 2 there are four parameters, indicating the expected number of drugs dispensed for each of the four values of the covariate.
@@
Table 3 displays the estimates for all of the model parameters. The narrow 95% credible intervals associated with these estimated parameters indicate low uncertainty within the estimated densities. The posterior median for the number of prescribed drugs rises by about a factor of four from state 1 to state 2, reflecting the worsening severity of patients’ health conditions. In state 2, expected drug dispensations for emergency department visits and hospitalizations are slightly higher than other types of visits. Point process rates are roughly 35% higher in State 2 than in State 1, suggesting increased healthcare system interactions during poorer health conditions. This indicates a correlation between disease severity and healthcare utilization, with state 2 indicating the higher level of healthcare engagement among patients. In the state transition dynamics, the median stay in State 1 is roughly $2.3$ years and the probability that a patient is still in State 1 after four years is approximately $0.30$. When a transition from state 1 occurs, the probability that it will be directly to the death state is $0.23$. The estimates for state 2 suggest a more persistent state (median stay being over 4 years) and potentially more severe state of illness. Patients in State 1 are less likely to die (or move from the province) in any given year, compared to those in State 2. The median time to death in state 2 is 4.3 years. We also notice that $\nu_{1}\approx\nu_{2}$ meaning that conditional on a healthcare visit, patients are almost equally likely to start in each of two states in the heathcare system.
@@
This inferred state dynamic provides valuable insights for healthcare management, offering guidance on the performance of current COPD interventions. The tendency for patients in state 1 to remain in that state initially before transitioning to state 2 underscores the need for interventions aimed at preventing or delaying disease progression, particularly during the early stages. For patients already in state 2, the observed persistence suggests a potentially more challenging treatment landscape, emphasizing the importance of comprehensive and long-term management strategies. Understanding these state transition dynamics informs healthcare providers and policymakers about the effectiveness of existing interventions and guides the development of targeted strategies to improve patient outcomes. By identifying critical transition points and patient populations at higher risk of disease progression, healthcare resources can be allocated more effectively to provide timely and appropriate interventions, ultimately improving the management and care of individuals with COPD. However, it is important to note that there is no guarantee that the latent state and trajectory groupings will directly correspond to the underlying biologic processes or disease stages as our initial model is built without clinical measurements. Therefore, the identified states correspond to the general health conditions of COPD patients, implying that a higher disease level or intensity of healthcare encounters does not strictly indicate a deterioration in COPD but could signify any temporary or prolonged period of poor health condition. By using Canadian cohort data, the initial latent groupings are expected to generate hypotheses about how the subjects in the dataset are being managed.
@@
$\lambda$ (Marked point process rate) 5.06 (4.97, 5.16) 6.82 (6.72, 6.93)

### v2

@@ Abstract
In medical research, understanding changes in outcome measurements is crucial for inferring shifts in health conditions. However, traditional methods often struggle with large, irregularly longitudinal data and fail to account for the tendency of individuals in poorer health to interact more frequently with the healthcare system. Additionally, clinical data can lack information on terminating events like death. To address these challenges, we start from the continuous- time hidden Markov model which models observed data as outcomes influenced by latent health states. Our extension incorporates a point process to account for the impact of health states on observation timings and includes a “death” state to model unobserved terminating events through a Poisson process, where transition rates depend on the latent health state. This approach captures both the severity of the disease and the timing of healthcare interactions. We present an exact Gibbs sampler procedure that alternates between sampling the latent health state paths and the model parameters. By including the “death” state, we mitigate biases in parameter estimation that would arise from solely modelling “live” health states. Simulation studies demonstrate that the proposed Gibbs sampler performs effectively. We apply our method to Canadian healthcare data, offering valuable insights for healthcare management.
@@ 1 Introduction
A multi-state model captures the status of an individual longitudinally as a time-discretization of a continuous-time Markov process, and in many settings the common assumption that the health status of a patient falls into one of a finite number of categories (usually ordered by severity) is a plausible one. Inference for continuous-time Markov processes observed discretely and either precisely or partially has been well studied (see, for example, Bladt and Sørensen, 2005; Pfeuffer, 2017). However, health system records typically do not include direct information about underlying health status; instead, it is inferred indirectly through a proxy measure. Such a process can be modelled as a continuous-time hidden Markov model (CTHMM, e.g. Cappé et al., 2005), which has been studied recently in medical applications such as disease progression (Luo et al., 2021) and cancer screening (Lange et al., 2015; Klausch et al., 2023). While a CTHMM can address the issue of non-equidistant longitudinal observations, in medical research, it is possible that the timing of healthcare utilization itself is informative in the assessment of health status progression. For example, patients who are in more severe health states typically interact with the healthcare (e.g. Cook and Lawless, 2018). However, if the intensity of the Poisson process depends on the state of a hidden Markov process, it is often referred to as a Markov-modulated Poisson process (e.g. Rydén, 1996; Fearnhead and Sherlock, 2006). There have been studies modelling this observation time process using frequentist approaches, such as Lange et al. (2015) and Mews et al. (2023).
In addition, data from the claim and health care system provide no information for terminating events, such as death. However, appropriate parameter inference will only follow if we distinguish a lack of interaction with the healthcare system due to, for example, mild illness from a lack of interaction due to the subject being dead. Nevala et al. (2024) included the death data from a different data system to model colorectal cancer progression. However, to the best of our knowledge, there is no existing method to account for the terminating event, such as death, when it is not available in a dataset and event times are informative. In this paper, we develop a Bayesian inference procedure that takes into account the observation process by modelling the observed data as an outcome process and the time of observations as a point process jointly alongside the CTHMM, given the underlying latent health state. Moreover, we also consider the death time as missing data, and model the death as an informative censoring. This extension allows us to model disease severity and death not only from the types of care received but also from the times between different observed events and their respective frequency. It also provides insights on how long patients tend to survive in each state.
@@
The paper is organized as follows. Section 2 demonstrate the model structure, followed by the likelihood construction in Section 2.1 and the forward-backward method to calculate the probability of each state at the observation time in Section 3. Section 4 layouts the Gibbs sampler to draw inference for the model, and some empirical studies, including simulation and real data analysis, are in Section 5. The paper concludes in Section 6 with a summary of findings and discusses areas for future research.
@@ 2 A Markov-modulated Poisson process with an outcome process
Conditional on the latent process being in state $k\in\mathcal{K}$, at time $\tau_{t}$, the outcome observation $o_{t}$ is drawn from a distribution with a probability mass or density function $f_{t}(o_{t}|X_{\tau_{t}}=k)\equiv f_{k}(o_{t}|z_{t},\beta_{k})$, where $z_{t}\in\mathbb{R}^{D}$ is a vector of covariates (including an intercept) and $\beta$ is a $D$-vector of covariate effects. In all of our simulations, including the real-data example, the prior for $\beta_{k}$, is chosen to be conditionally conjugate to $f_{k}$ so that it is straightforward to sample from the conditional posterior for each $\beta_{k}$ given $X_{\tau_{1}},\dots,X_{\tau_{T}}$. It is possible to extend inference to the non-conjugate case using the Metropolis–Hastings algorithm; see the supplementary material for a brief discussion. Define the matrix $\mathbf{B}=(\beta_{d,k})$ for $d=1,\ldots,D$ and $k=1,\ldots,K$ as the regression coefficient matrix.
@@ 2.1 Likelihood for the MMPP with an outcome process
$\mathcal{L}(\mathbf{Q},\nu|x_{[0,\tau]})=f(x_{[0,\tau]}|\mathbf{Q},\nu)=\nu_{x_{0}}\prod\limits_{l=1}^{K}{\prod\limits_{m\neq l}{{q_{l,m}}^{{N_{l,m}}}\exp\left({-{q_{l,m}}{R_{l}}}\right)}},$
@@
If independent Gamma-distributed priors are assigned to all of the components $\lambda_{i}$ and $q_{l,m}$ (see Section 4.1) then, conditional on the complete chains for all individuals, the posteriors for these parameters also have independent Gamma distributions. This permits the use of the Gibbs steps described in Section 4.2. An alternative partial imputation, which necessitates Metropolis–Hastings updates for these parameters, is discussed in the Supplementary Material.
@@ 3 The forward-backward algorithm
$\left(\begin{array}[]{cc}Q-\mathbf{\Lambda}&\boldsymbol{\lambda}^{\top}\\ \mathbf{0}^{\top}&0\\ \end{array}\right)$
where $\mathbf{\Lambda}=\text{diag}\left(\lambda_{1},\ldots,\lambda_{K}\right)$. Writing $Y_{\tau^{-}}$ for $\lim_{s\uparrow\tau}Y_{s}$, and defining $\eta_{t-1,t,i,k}:=\left[\exp\left\{(Q-\mathbf{\Lambda})(\tau_{t}-\tau_{t-1})\right\}\right]_{i,k}$, we see that
$\mathbb{P}\left(\tau_{t},X_{\tau_{t}}=k\left|X_{{\tau_{t-1}}}=i\right.\right)=\mathbb{P}\left(Y_{\tau_{t}^{-}}=k\left|Y_{\tau_{t-1}}=i\right.\right)\lambda_{k}=\left[\exp\left\{(Q-\mathbf{\Lambda})(\tau_{t}-\tau_{t-1})\right\}\right]_{i,k}\lambda_{k}=\eta_{t-1,t,i,k}\lambda_{k}.$
@@
$\mathbb{P}(X_{\tau_{e}}=k,\tau_{e}|X_{\tau_{T}}=i)=\left[\exp\left\{(Q-\mathbf{\Lambda})(\tau_{e}-\tau_{T})\right\}\right]_{i,k}=:\eta_{T,e,i,k}.$
@@
From its definition, $a_{e,k}\propto\mathbb{P}(X_{\tau_{e}}=k|o_{1:T},\tau_{1:T},\tau_{e})$ so we may sample a value for $x_{\tau_{e}}$ with the probabilities of the individual states proportional to the elements of $a_{e}$. Suppose that we have simulated $x_{\tau_{e}},x_{\tau_{T}},\dots,x_{\tau_{t+1}}$ from their joint distribution given $o_{1:T}$, $\tau_{1:T}$ and $\tau_{e}$. We now describe how to simulate $x_{\tau_{t}}$ backward from its conditional distribution given $x_{\tau_{e}},x_{\tau_{T}},\dots,x_{\tau_{t+1}}$, $o_{1:T}$, $\tau_{1:T}$ and $\tau_{e}$. From the hidden-Markov structure, conditional on $x_{\tau_{t+1}}$, all random variables with time indices before $t+1$ are independent of those with indices after $t+1$, so
@@ 5.1 Example 1
For each of $N=50$ subjects, we simulate a realization of a two-state latent process without the death, $X_{[0,5]}$, using a generator of
@@
Trace plots, auto-correlations plots and histograms for $\lambda_{1}$ and $q_{12}$ are shown in Figures 3 and 4; Integrated auto correlation times (IACTs) for $\lambda_{1}$, $\lambda_{2}$, $q_{1,2}$, $q_{2,1}$, $\beta_{1}$ and $\beta_{2}$ are provided in Table 1. In Scenario I, where the data are generated from a model characterized by clear separation between the states, the Gibbs sampler performs well, with fast convergence and mixing. In Scenario II, where the timing of the events provides no information about the state, the Gibbs sampler’s performance declines slightly, though it still remains promising. The posterior distribution of $\lambda_{1}$ is slightly off the true value. Challenges arise when the data are generated from a less well-separated outcome process, leading to notable auto-correlations in the posterior samples, particularly those associated with $\mathbf{Q}$. The reduction in mixing efficiency is considerable; however, the chain does still mix adequately over the 20,000 iterations shown. By contrast, in Scenario IV, when the point-process information is removed from Scenario III, the exploration of the posterior over the 20,000 iterations is so poor that we would judge that the chain has perhaps not yet even converged. Moreover, the posterior for $q_{1,2}$ includes values that are not strongly supported by the data. This illustrates the most important point: especially in cases where the outcome process is not particularly informative, the information from the timings of the events provides an important contribution to the posterior and, possibly because of the tighter posterior and more clearly defined paths for each $X_{[0,\tau]}$, improves the mixing of the chain.
@@ 5.2 Example 2
In this example, we demonstrate valid inference for simulated data with a similar pattern to our real data. We generate the data from a three-state model which is similar to that used for our real data, with $N=500$ subjects. For each subject, we simulate a realization of a three-state latent process with the death as the absorbing state, $X_{[0,10]}$,
@@
This is equivalent to $O_{t}|(Z_{t}=1)\sim\mathsf{Poisson}(\mu_{1,k})$ and $O_{t}|(Z_{t}=0)\sim\mathsf{Poisson}(\mu_{0,k})$, with $\log\mu_{1,k}=B_{1,k}+B_{2,k}$ and $\log\mu_{0,k}=B_{1,k}$. In this case, the outcome expectations (marginalized over the covariate) are $0.35\mu_{0,k}+0.65\mu_{1,k}$, which are $0.46$ and $1.71$ for the two states, respectively. The point process rate is set as $\lambda=(6,10,0)^{\top}$. State 3 is the death state without any observed outcome. If the patient progresses to the absorbing state before the observation window, i.e. death, then the last observation is the final one before death. Otherwise, the last observation is the one right before the observation window, $\tau=10$. We assumed that the first observation is made at time 0 according to the initial distribution. Similar to Example 1, we place non-informative priors on all the parameters in this study. Each of $\lambda_{1}$, $\lambda_{2}$, $q_{12}$ and $q_{21}$ have independent $\mathsf{Gamma}(1,1/8)$ priors. We place independent $\mathsf{Gamma}(0.1,0.1)$ priors on each of the four Poisson expectations, $\mu_{i,k}$, $i=0,1$, $k=1,2$, and a $\mathsf{Beta}(1,1)$ prior for $\nu_{1}$.
Table 2 presents the integrated auto correlation times (IACTs) for $\boldsymbol{\lambda}$, $\mathbf{Q}$ and $\mathbf{B}$ over 20,000 iterations and shows that all of these parameters mix well. Figure 5 presents a histogram of the marginal posterior for each of these parameters, with each true parameter value falling within the main posterior mass. This example demonstrate that by simulating the data-induced window similar to the real data, we can achieve valid inference.
@@ 5.3 Application: Canadian healthcare system data
In this section, we delve into the proposed methodology designed to enhance our understanding of the progression of chronic diseases, particularly COPD. The significance of this understanding lies in its potential to help with effective management of healthcare systems and pave the way for methods of early diagnosis and personalized care, as the model can help the management team understand the “care path” or “trajectory” of the cohort before they become severely ill. While clinical and administrative data harbor valuable insights, traditional methods for modelling disease progression face limitations when dealing with data collected at irregular intervals and interaction frequency, such as when patients choose to engage with the healthcare system.
@@
Utilizing established case definitions based on diagnostic codes (Lix et al., 2018), a total of 76,888 COPD patients were enrolled, with an incident event (ICD-9 491x, 492x, 496x; ICD-10 J41-J44) occurring after a minimum of two years at risk with no events. The observation period spanned from January 1998, commencing with the patients’ first diagnosis, until December 2014. Physicians observed these patients solely during medical visits, which transpired when patients chose to interact with the healthcare system. In these medical visits, pertinent medical information was gathered, and from this observed information, we aim to infer the patients’ unobserved disease status, which is represented using a discrete disease state model. In this study, the disease status was indirectly gauged through a proxy: the number of prescribed medications at the time of observation. Notably, this information was available exclusively for patients with drug insurance, leading us to focus on patients aged over 65 years, constituting the study population for our analysis. This full dataset has been previously analyzed in Luo et al. (2021). Due to the data restriction, we are allowed to randomly sample 1,000 patients to demonstrate our proposed method. Figure 6 shows one COPD patient’s follow up from this dataset. As shown in the figure, we observe the number of drugs prescribed as the observation process. In addition, we have also observe variations in the type of encounters, i.e., general practitioner visit (GP), hospitalization (HOSP), specialist visit (SPEC) and emergency department visit (ED), which will be incorporated as time-varying covariates in our model. These changes in the observation point process, transitioning from sparse to dense, reflect the evolving severity of the chronic disease under study. In the dataset, the average number of drugs prescribed from a visit for GP, ED, HOSP, and SPEC is 5.06, 5.57, 5.28, and 5.17 respectively, suggesting little deviation among the types of healthcare utilization. Over the 1,000 patients, the average time from the first recorded interaction until the last observation is about 13.71 years.
@@
We apply the model from Section 2 to this sample dataset, with the inference methodology from Section 4. We focus on a three-state model with State 3 (the unobserved death state) as an absorbing state. We place independent $\mathsf{Gamma}(1,1/8)$ priors on each $\lambda_{i}$ and all elements in $\mathbf{Q}$, and a $\mathsf{Gamma}(0.1,0.1)$ prior on each expected number of drugs dispensed for each healthcare utilization, and a $\mathsf{Beta}(1,1)$ prior for $(\nu_{1},\nu_{2})$, with $\nu_{3}=0$. The outcome process is modeled with a Poisson distribution, with the types of health care utilization (HOSP, SPEC, GP, and ED) as a time-dependent covariate; equivalently, for each of states 1 and 2 there are four parameters, indicating the expected number of drugs dispensed for each of the four values of the covariate.
@@
We ran the algorithm for 20,000 iterations; for all parameters, the estimated IACTs were less than 5, indicating excellent mixing of the MCMC algorithm. Table 3 displays the estimates for all of the model parameters. The narrow 95% credible intervals associated with these estimated parameters indicate low uncertainty within the estimated densities. The posterior median for the number of prescribed drugs rises by about a factor of four from State 1 to State 2, reflecting the worsening severity of patients’ health conditions. In State 2, expected drug dispensations for emergency department visits and hospitalizations are slightly higher than other types of visits. Point process rates are roughly 35% higher in State 2 than in State 1, suggesting increased healthcare system interactions during poorer health conditions. This indicates a correlation between disease severity and healthcare utilization, with state 2 indicating the higher level of healthcare engagement among patients. In the state transition dynamics, the median stay in State 1 is roughly $2.3$ years and the probability that a patient is still in State 1 after four years is approximately $0.30$. When a transition from state 1 occurs, the probability that it will be directly to the death state is $0.23$. The estimates for state 2 suggest a more persistent state (median stay being over 4 years) and potentially more severe state of illness. Patients in State 1 are less likely to die (or move from the province) in any given year, compared to those in State 2. The median time to death in State 2 is 4.3 years. We also notice that $\nu_{1}\approx\nu_{2}$ meaning that conditional on a healthcare visit, patients are almost equally likely to start in each of the two states in the heathcare system.
We also ran the Gibbs sampler without the point process as with Scenario IV in Example 1. The parameter estimates are reproduced in the Supplementary Material. They show that ignoring the point process information makes patients seem to remain in State 1 for longer. The other substantial difference is that the transition rate to the death state from State 1 is slightly greater than that from State 2. We also fitted a four-state model, again with only transitions to higher states allowed, and now with State 4 (the unobserved death state) as an absorbing state. We compared a Laplace-based approximation of its marginal likelihood with a similar approximation to the marginal likelihood of the three-state model; see the Supplementary Material for details. The estimated marginal likelihoods for three-state and four-state models are $-2702.93$ and $-2687.05$ respectively, leading to corresponding BICs of $5502.57,5612.19$. This shows that the three-state model is preferable; results for the four-state model are also presented in the Supplementary Material.
@@
This inferred state dynamic provides valuable insights for healthcare management, offering guidance on the performance of current COPD interventions. The tendency for patients in State 1 to remain in that state initially before transitioning to State 2 underscores the need for interventions aimed at preventing or delaying disease progression, particularly during the early stages. For patients already in State 2, the observed persistence suggests a potentially more challenging treatment landscape, emphasizing the importance of comprehensive and long-term management strategies. Understanding these state transition dynamics informs healthcare providers and policymakers about the effectiveness of existing interventions and guides the development of targeted strategies to improve patient outcomes. By identifying critical transition points and patient populations at higher risk of disease progression, healthcare resources can be allocated more effectively to provide timely and appropriate interventions, ultimately improving the management and care of individuals with COPD. However, it is important to note that there is no guarantee that the latent state and trajectory groupings will directly correspond to the underlying biologic processes or disease stages as our initial model is built without clinical measurements. Therefore, the identified states correspond to the general health conditions of COPD patients, implying that a higher disease level or intensity of healthcare encounters does not strictly indicate a deterioration in COPD but could signify any temporary or prolonged period of poor health condition. By using this Canadian cohort dataset, the initial latent groupings are expected to generate hypotheses about how the patients are being managed.
@@
$\lambda$ (Marked point process rate) 5.07 (4.97, 5.16) 6.82 (6.72, 6.93)